# SVDQuant: Absorbing Outliers by Low-Rank Components for 4-Bit Diffusion Models

### v3

@@ 1 Introduction
Diffusion models have shown remarkable capabilities in generating high-quality images (Ho et al., 2020), with recent advances further enhancing user control over the generation process. Trained on vast data, these models can create stunning images from simple text prompts, unlocking diverse image editing and synthesis applications (Meng et al., 2022b; Ruiz et al., 2023; Zhang et al., 2023).
@@
To pursue higher image quality and more precise text-to-image alignment, researchers are scaling up diffusion models. As shown in Figure 2, Stable Diffusion (SD) (Rombach et al., 2022) 1.4 only has 800M parameters, while SDXL (Podell et al., 2024) scales this up to 2.6B parameters. AuraFlow v0.1 (fal.ai, 2024) extends this further to 6B parameters, with the latest model, FLUX.1 (Black-Forest-Labs, 2024), pushing the boundary to 12B parameters. Compared to large language models (LLMs), diffusion models are significantly more computationally intensive. Their computational costsMeasured by the number of Multiply-Accumulate operations (MACs). 1 MAC=2 FLOPs. increase more rapidly with model size, posing a prohibitive memory and latency barrier for real-world model deployment, particularly for interactive use cases that demand low latency.
@@
As Moore’s law slows down, hardware vendors are turning to low-precision inference to sustain performance improvements. For instance, NVIDIA’s Blackwell Tensor Cores introduce a new 4-bit floating point (FP4) precision, doubling the performance compared to FP8 (NVIDIA, 2024). Therefore, using 4-bit inference to accelerate diffusion models is appealing. In the realm of LLMs, researchers have leveraged quantization to compress model sizes and boost inference speed (Dettmers et al., 2022; Xiao et al., 2023). However, unlike LLMs–where latency is primarily constrained by loading model weights on modern GPUs, especially with small batch sizes–diffusion models are heavily computationally bounded, even with a single batch. As a result, weight-only quantization cannot accelerate diffusion models. To achieve speedup on these devices, both weights and activations must be quantized to the same bit width; otherwise, the lower-precision weight will be upcast during computation, negating potential performance enhancements.
@@
In this work, we focus on quantizing both the weights and activations of diffusion models to 4 bits. This challenging and aggressive scheme is often prone to severe quality degradation. Existing methods like smoothing (Xiao et al., 2023; Lin et al., 2024), which transfer the outliers between the weights and activations, are less effective since both sides are highly vulnerable to outliers. To address this issue, we propose a general-purpose quantization paradigm, SVDQuant. Our core idea is to use a low-cost branch to absorb outliers on both sides. To achieve this, as illustrated in Figure 3, we first aggregate the outliers by migrating them from activation ${\bm{X}}$ to weight ${\bm{W}}$ via smoothing. Then we apply Singular Value Decomposition (SVD) to the updated weight, $\hat{{\bm{W}}}$, decomposing it into a low-rank branch ${\bm{L}}_{1}{\bm{L}}_{2}$ and a residual $\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}$. The low-rank branch operates at 16 bits, allowing us to quantize only the residual to 4 bits, significantly reducing outliers and magnitude. However, naively running the low-rank branch separately incurs substantial memory access overhead, offsetting the speedup of 4-bit inference. To overcome this, we co-design a specialized inference engine Nunchaku, which fuses the low-rank branch computation into the 4-bit quantization and computation kernels. This design enables us to achieve measured inference speedup even with additional branches.
@@
SVDQuant can quantize various text-to-image diffusion architectures into 4 bits, including both UNet (Ho et al., 2020; Ronneberger et al., 2015) and DiT (Peebles & Xie, 2023) backbones, while maintaining visual quality. It supports both INT4 and FP4 data types and integrates seamlessly with pre-trained low-rank adapters (LoRA) (Hsu et al., 2022) without requiring re-quantization. To our knowledge, we are the first to successfully apply 4-bit post-training quantization to both the weights and activations of diffusion models, and achieve measured speedup on NVIDIA GPUs. On the latest 12B FLUX.1, our 4-bit models largely preserve the image quality and reduce the memory footprint of the original BF16 model by 3.5×. Furthermore, our INT4 and FP4 model delivers a 3.0× and 3.1× speedup over the NF4 weight-only quantized baseline on the 16GB laptop-level RTX 4090 and desktop-level RTX 5090 GPU, respectively. See Figure 1 for visual examples.
@@ 2 Related Work
Diffusion models  (Sohl-Dickstein et al., 2015; Ho et al., 2020) have emerged as a powerful class of generative models, known for generating high-quality samples by modeling the data distribution through an iterative denoising process. Recent advancements in text-to-image diffusion models (Balaji et al., 2022; Rombach et al., 2022; Podell et al., 2024) have already revolutionized content generation. Researchers further shifted from convolution-based UNet architectures (Ronneberger et al., 2015; Ho et al., 2020) to transformers (Peebles & Xie, 2023; Bao et al., 2023) and scaled up the model size (Esser et al., 2024). However, diffusion models suffer from extremely slow inference speed due to their long denoising sequences and intense computation. To address this, various approaches have been proposed, including few-step samplers (Zhang & Chen, 2022; Zhang et al., 2022; Lu et al., 2022) or distilling fewer-step models from pre-trained ones (Salimans & Ho, 2021; Meng et al., 2022a; Song et al., 2023; Luo et al., 2023; Sauer et al., 2024; Yin et al., 2024b; a; kan, 2024). Another line of works choose to optimize or accelerate computation via efficient architecture design (Li et al., 2023b; 2020; Cai et al., 2024; Liu et al., 2024a), quantization (Shang et al., 2023; Li et al., 2023a), sparse inference (Li et al., 2022; Ma et al., 2024c; b), and distributed inference (Li et al., 2024b; Wang et al., 2024c; Chen et al., 2024b). This work focuses on quantizing the diffusion models to 4 bits to reduce the computation complexity. Our method can also be applied to few-step diffusion models to further reduce the latency (see Section 5.2).
@@
Quantization has been recognized as an effective approach for LLMs to reduce the model size and accelerate inference (Dettmers et al., 2022; Frantar et al., 2023; Xiao et al., 2023; Lin et al., 2025; 2024; Kim et al., 2024; Zhao et al., 2024d). For diffusion models, Q-Diffusion (Li et al., 2023a) and PTQ4DM (Shang et al., 2023) first achieved 8-bit quantization. Subsequent works refined these techniques with approaches like sensitivity analysis (Yang et al., 2023) and timestep-aware quantization (He et al., 2023; Huang et al., 2024; Liu et al., 2024b; Wang et al., 2024a). Some recent works extended the setting to text-to-image models (Tang et al., 2024; Zhao et al., 2024c), DiT backbones (Wu et al., 2024), quantization-aware training (He et al., 2024; Zheng et al., 2024; Wang et al., 2024b; Sui et al., 2024), video generation (Zhao et al., 2024b), and different data types (Liu & Zhang, 2024). Among these works, only MixDQ (Zhao et al., 2024c) and ViDiT-Q (Zhao et al., 2024b) implement low-bit inference engines and report measured 8-bit speedup on GPUs. In this work, we push the boundary further by quantizing ffusion models to 4 bits, supporting both the integer or floating-point data types, compatible with the UNet backbone (Ho et al., 2020) and recent DiT architecture (Peebles & Xie, 2023). Our co-designed inference engine, Nunchaku, further ensures on-hardware speedup. Additionally, when applying LoRA to the model, existing methods require fusing the LoRA branch to the main branch and re-quantizing the model to avoid tremendous memory-access overhead in the LoRA branch. Nunchaku cuts off this overhead via kernel fusion, allowing the low-rank branch to run efficiently as a separate branch, eliminating the need for re-quantization.
@@
Low-rank decomposition has gained significant attention in deep learning for enhancing computational and memory efficiency (Hu et al., 2022; Zhao et al., 2024a; Jaiswal et al., 2024). While directly applying this approach to model weights can reduce the compute and memory demands (Hsu et al., 2022; Yuan et al., 2023; Li et al., 2023c), it often leads to performance degradation. Instead, Yao et al. (2024) combined it with quantization for model compression, employing a low-rank branch to compensate for the quantization error. Low-Rank Adaptation (LoRA) (Hu et al., 2022) introduces another active line of research using low-rank matrices to adjust a subset of pre-trained weights for efficient fine-tuning. This has sparked numerous advancements (Dettmers et al., 2023; Guo et al., 2024; Li et al., 2024c; Xu et al., 2024b; Meng et al., 2024), which combines quantized models with low-rank adapters to reduce memory usage during model fine-tuning. However, our work differs in two major aspects. First, our goal is different, as we aim to accelerate model inference through quantization, while previous works focus on model compression or efficient fine-tuning. Thus, they primarily consider weight-only quantization, resulting in no speedup. Second, as shown in our experiments (Figure 6 and ablation study in Section 5.2), directly applying these methods not only degrades the image quality, but also introduces significant overhead. In contrast, our method yields much better performance due to our joint quantization of weights and activations and overhead reduction of our inference engine Nunchaku.
@@ 3 Quantization Preliminary
[…]
In this work, we focus on W4A4 quantization for acceleration, where outliers in both weights and activations place substantial obstacles. Traditional methods to suppress these outliers include quantization-aware training (QAT) (He et al., 2024) and rotation (Ashkboos et al., 2024; Liu et al., 2024c; Lin et al., 2025). QAT requires massive computing resources, especially for tuning models with more than 10B parameters such as FLUX.1. Rotation is inapplicable due to the usage of adaptive normalization layers (Peebles & Xie, 2023) in diffusion models. The runtime-generated normalization weights preclude the offline rotation with the weights of projection layers, while online rotation of both activations and weights incurs significant runtime overhead.
@@ 4 Method
Figure 4: Example value distribution of inputs and weights in PixArt-$\Sigma$ (Chen et al., 2024a) ${\bm{\lambda}}$ is the smooth factor. Red indicates the outliers. Initially, both the input ${\bm{X}}$ and weight ${\bm{W}}$ contain significant outliers. After smoothing, the range of $\hat{{\bm{X}}}$ is reduced with much fewer outliers, while $\hat{{\bm{W}}}$ shows more outliers. Once the SVD low-rank branch ${\bm{L}}_{1}{\bm{L}}_{2}$ is subtracted, the residual ${\bm{R}}$ has a narrower range and is free from outliers.
@@ 4.2 SVDQuant: Absorbing Outliers via Low-Rank Branch
Migrate outliers from activation to weight. Smoothing (Xiao et al., 2023; Lin et al., 2024) is an effective approach for reducing outliers. We can smooth outliers in activations by scaling down the input ${\bm{X}}$ and adjusting the weight matrix ${\bm{W}}$ correspondingly using a per-channel smoothing factor ${\bm{\lambda}}\in{\mathbb{R}}^{m}$. As shown in Figure 4(a)(c), the smoothed input $\hat{{\bm{X}}}={\bm{X}}\cdot\text{diag}({\bm{\lambda}})^{-1}$ exhibits reduced magnitude and fewer outliers, resulting in lower input quantization error. However, in Figure 4(b)(d), the transformed weight $\hat{{\bm{W}}}={\bm{W}}\cdot\text{diag}({\bm{\lambda}})$ has a significant increase in both magnitude and the presence of outliers, which in turn raises the weight quantization error. Consequently, the overall error reduction is limited.
@@
See Appendix 4.2 for the proof. From this proposition, we obtain the intuition that the quantization error $\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}$ is bounded by the magnitude of the residual $\left\lVert{\bm{R}}\right\rVert_{F}$. Thus, our goal is to find the optimal ${\bm{L}}_{1}{\bm{L}}_{2}$ that minimizes $\left\lVert{\bm{R}}\right\rVert_{F}=\left\lVert\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}\right\rVert_{F}$, which can be solved by Singular Value Decomposition (SVD) (Eckart & Young, 1936; Mirsky, 1960). Given the SVD of $\hat{{\bm{W}}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}$, the optimal solution is ${\bm{L}}_{1}={\bm{U}}{\bm{\Sigma}}_{:,:r}$ and ${\bm{L}}_{2}={\bm{V}}_{:r,:}$. Figure 5 illustrates the singular value distribution of the original weight ${\bm{W}}$, transformed weight $\hat{{\bm{W}}}$ and residual ${\bm{R}}$. The singular values of the original weight ${\bm{W}}$ are highly imbalanced. After smoothing, the singular value distribution of $\hat{{\bm{W}}}$ becomes even sharper, with only the first several values being significantly larger. By removing these dominant values, the magnitude of the residual ${\bm{R}}$ is dramatically reduced, as $\left\lVert{\bm{R}}\right\rVert_{F}=\sqrt{\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}}$, compared to the original magnitude $\left\lVert\hat{{\bm{W}}}\right\rVert_{F}=\sqrt{\sum_{i=1}^{\min(m,n)}\sigma_{i}^{2}}$, where $\sigma_{i}$ is the $i$-th singular value of $\hat{{\bm{W}}}$. Furthermore, Figure 4(d)(e) show that ${\bm{R}}$ exhibits fewer outliers with a substantially compressed value range compared to $\hat{{\bm{W}}}$. In practice, we further reduce quantization errors by iteratively updating the low-rank branch through decomposing ${\bm{W}}-Q({\bm{R}})$ and adjusting ${\bm{R}}$ accordingly for several iterations, and then picking the result with the smallest error.

### v4

@@ 1 Introduction
Diffusion models have shown remarkable capabilities in generating high-quality images (ho2020denoising), with recent advances further enhancing user control over the generation process. Trained on vast data, these models can create stunning images from simple text prompts, unlocking diverse image editing and synthesis applications (meng2022sdedit; ruiz2023dreambooth; zhang2023adding).
@@
To pursue higher image quality and more precise text-to-image alignment, researchers are scaling up diffusion models. As shown in Figure 2, Stable Diffusion (SD) (rombach2022high) 1.4 only has 800M parameters, while SDXL (podell2023sdxl) scales this up to 2.6B parameters. AuraFlow v0.1 (auraflow0.1) extends this further to 6B parameters, with the latest model, FLUX.1 (flux1), pushing the boundary to 12B parameters. Compared to large language models (LLMs), diffusion models are significantly more computationally intensive. Their computational costsMeasured by the number of Multiply-Accumulate operations (MACs). 1 MAC=2 FLOPs. increase more rapidly with model size, posing a prohibitive memory and latency barrier for real-world model deployment, particularly for interactive use cases that demand low latency.
@@
As Moore’s law slows down, hardware vendors are turning to low-precision inference to sustain performance improvements. For instance, NVIDIA’s Blackwell Tensor Cores introduce a new 4-bit floating point (FP4) precision, doubling the performance compared to FP8 (blackwell). Therefore, using 4-bit inference to accelerate diffusion models is appealing. In the realm of LLMs, researchers have leveraged quantization to compress model sizes and boost inference speed (dettmers2022gpt3; xiao2023smoothquant). However, unlike LLMs–where latency is primarily constrained by loading model weights on modern GPUs, especially with small batch sizes–diffusion models are heavily computationally bounded, even with a single batch. As a result, weight-only quantization cannot accelerate diffusion models. To achieve speedup on these devices, both weights and activations must be quantized to the same bit width; otherwise, the lower-precision weight will be upcast during computation, negating potential performance enhancements.
@@
In this work, we focus on quantizing both the weights and activations of diffusion models to 4 bits. This challenging and aggressive scheme is often prone to severe quality degradation. Existing methods like smoothing (xiao2023smoothquant; lin2024awq), which transfer the outliers between the weights and activations, are less effective since both sides are highly vulnerable to outliers. To address this issue, we propose a general-purpose quantization paradigm, SVDQuant. Our core idea is to use a low-cost branch to absorb outliers on both sides. To achieve this, as illustrated in Figure 3, we first aggregate the outliers by migrating them from activation $\bm{X}$} to weight $\bm{W}$} via smoothing. Then we apply Singular Value Decomposition (SVD) to the updated weight, $\hat{{\bm{W}}}$, decomposing it into a low-rank branch ${\bm{L}}_{1}{\bm{L}}_{2}$ and a residual $\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}$. The low-rank branch operates at 16 bits, allowing us to quantize only the residual to 4 bits, significantly reducing outliers and magnitude. However, naively running the low-rank branch separately incurs substantial memory access overhead, offsetting the speedup of 4-bit inference. To overcome this, we co-design a specialized inference engine Nunchaku, which fuses the low-rank branch computation into the 4-bit quantization and computation kernels. This design enables us to achieve measured inference speedup even with additional branches.
@@
SVDQuant can quantize various text-to-image diffusion architectures into 4 bits, including both UNet (ho2020denoising; ronneberger2015u) and DiT (peebles2023scalable) backbones, while maintaining visual quality. It supports both INT4 and FP4 data types and integrates seamlessly with pre-trained low-rank adapters (LoRA) (hsu2022lowrank) without requiring re-quantization. To our knowledge, we are the first to successfully apply 4-bit post-training quantization to both the weights and activations of diffusion models, and achieve measured speedup on NVIDIA GPUs. On the latest 12B FLUX.1, our 4-bit models largely preserve the image quality and reduce the memory footprint of the original BF16 model by 3.5×. Furthermore, our INT4 and FP4 model delivers a 3.0× and 3.1× speedup over the NF4 weight-only quantized baseline on the 16GB laptop-level RTX 4090 and desktop-level RTX 5090 GPU, respectively. See Figure 1 for visual examples.
@@ 2 Related Work
Diffusion models  (sohl2015deep; ho2020denoising) have emerged as a powerful class of generative models, known for generating high-quality samples by modeling the data distribution through an iterative denoising process. Recent advancements in text-to-image diffusion models (balaji2022ediffi; rombach2022high; podell2023sdxl) have already revolutionized content generation. Researchers further shifted from convolution-based UNet architectures (ronneberger2015u; ho2020denoising) to transformers (peebles2023scalable; bao2023all) and scaled up the model size (esser2024scaling). However, diffusion models suffer from extremely slow inference speed due to their long denoising sequences and intense computation. To address this, various approaches have been proposed, including few-step samplers (zhang2022fast; zhang2022gddim; lu2022dpm) or distilling fewer-step models from pre-trained ones (salimans2021progressive; meng2022distillation; song2023consistency; luo2023latent; sauer2023adversarial; yin2024one; yin2024improved; kang2024distilling). Another line of works choose to optimize or accelerate computation via efficient architecture design (li2023snapfusion; li2020gan; cai2024condition; liu2024linfusion), quantization (shang2023post; li2023q), sparse inference (li2022efficient; ma2024deepcache; ma2024learning), and distributed inference (li2024distrifusion; wang2024pipefusion; chen2024asyncdiff). This work focuses on quantizing the diffusion models to 4 bits to reduce the computation complexity. Our method can also be applied to few-step diffusion models to further reduce the latency (see Section 5.2).
@@
Quantization has been recognized as an effective approach for LLMs to reduce the model size and accelerate inference (dettmers2022gpt3; frantar-gptq; xiao2023smoothquant; lin2025qserve; lin2024awq; kim2024squeezellm; zhao2024atom). For diffusion models, Q-Diffusion (li2023q) and PTQ4DM (shang2023post) first achieved 8-bit quantization. Subsequent works refined these techniques with approaches like sensitivity analysis (yang2023efficient) and timestep-aware quantization (he2023ptqd; huang2024tfmq; liu2024enhanced; wang2024towards). Some recent works extended the setting to text-to-image models (tang2023post; zhao2024mixdq), DiT backbones (wu2024ptq4dit), quantization-aware training (heefficientdm; zheng2024binarydm; wang2024quest; sui2024bitsfusion), video generation (zhao2024vidit), and different data types (liu2024hq). Among these works, only MixDQ (zhao2024mixdq) and ViDiT-Q (zhao2024vidit) implement low-bit inference engines and report measured 8-bit speedup on GPUs. In this work, we push the boundary further by quantizing ffusion models to 4 bits, supporting both the integer or floating-point data types, compatible with the UNet backbone (ho2020denoising) and recent DiT architecture (peebles2023scalable). Our co-designed inference engine, Nunchaku, further ensures on-hardware speedup. Additionally, when applying LoRA to the model, existing methods require fusing the LoRA branch to the main branch and re-quantizing the model to avoid tremendous memory-access overhead in the LoRA branch. Nunchaku cuts off this overhead via kernel fusion, allowing the low-rank branch to run efficiently as a separate branch, eliminating the need for re-quantization.
@@
Low-rank decomposition has gained significant attention in deep learning for enhancing computational and memory efficiency (lora; zhao2024galore; jaiswal2024welore). While directly applying this approach to model weights can reduce the compute and memory demands (hsu2022lowrank; yuan2023asvd; li2023losparse), it often leads to performance degradation. Instead, yao2023zeroquant combined it with quantization for model compression, employing a low-rank branch to compensate for the quantization error. Low-Rank Adaptation (LoRA) (lora) introduces another active line of research using low-rank matrices to adjust a subset of pre-trained weights for efficient fine-tuning. This has sparked numerous advancements (dettmers2023qlora; guo2024lq; li2024loftq; xu2024qalora; meng2024pissa), which combines quantized models with low-rank adapters to reduce memory usage during model fine-tuning. However, our work differs in two major aspects. First, our goal is different, as we aim to accelerate model inference through quantization, while previous works focus on model compression or efficient fine-tuning. Thus, they primarily consider weight-only quantization, resulting in no speedup. Second, as shown in our experiments (Figure 6 and ablation study in Section 5.2), directly applying these methods not only degrades the image quality, but also introduces significant overhead. In contrast, our method yields much better performance due to our joint quantization of weights and activations and overhead reduction of our inference engine Nunchaku.
@@ 3 Quantization Preliminary
The same approximation applies to convolutional layers. To speed up computation, modern commercial GPUs require both ${\bm{Q}}_{\bm{X}}$ and ${\bm{Q}}_{\bm{W}}$ using the same bit width. Otherwise, the low-bit weights need to be upcast to match the higher bit width of activations, or vice versa, negating the speed advantage. Following the notation in QServe (lin2025qserve), we denote $x$-bit weight, $y$-bit activation as WxxAyy. “INT” and “FP” refer to the integer and floating-point data types, respectively.
In this work, we focus on W4A4 quantization for acceleration, where outliers in both weights and activations place substantial obstacles. Traditional methods to suppress these outliers include quantization-aware training (QAT) (heefficientdm) and rotation (ashkboos2024quarot; liu2024spinquant; lin2025qserve). QAT requires massive computing resources, especially for tuning models with more than 10B parameters such as FLUX.1. Rotation is inapplicable due to the usage of adaptive normalization layers (peebles2023scalable) in diffusion models. The runtime-generated normalization weights preclude the offline rotation with the weights of projection layers, while online rotation of both activations and weights incurs significant runtime overhead.
@@ 4 Method
Figure 4: Example value distribution of inputs and weights in PixArt-$\Sigma$ (chen2024pixart) $\bm{\lambda}$} is the smooth factor. Red indicates the outliers. Initially, both the input $\bm{X}$} and weight $\bm{W}$} contain significant outliers. After smoothing, the range of $\hat{{\bm{X}}}$ is reduced with much fewer outliers, while $\hat{{\bm{W}}}$ shows more outliers. Once the SVD low-rank branch ${\bm{L}}_{1}{\bm{L}}_{2}$ is subtracted, the residual $\bm{R}$} has a narrower range and is free from outliers.
@@ 4.2 SVDQuant: Absorbing Outliers via Low-Rank Branch
Migrate outliers from activation to weight. Smoothing (xiao2023smoothquant; lin2024awq) is an effective approach for reducing outliers. We can smooth outliers in activations by scaling down the input $\bm{X}$} and adjusting the weight matrix $\bm{W}$} correspondingly using a per-channel smoothing factor ${\bm{\lambda}}\in{\mathbb{R}}^{m}$. As shown in Figure 4(a)(c), the smoothed input $\hat{{\bm{X}}}={\bm{X}}\cdot\text{diag}({\bm{\lambda}})^{-1}$ exhibits reduced magnitude and fewer outliers, resulting in lower input quantization error. However, in Figure 4(b)(d), the transformed weight $\hat{{\bm{W}}}={\bm{W}}\cdot\text{diag}({\bm{\lambda}})$ has a significant increase in both magnitude and the presence of outliers, which in turn raises the weight quantization error. Consequently, the overall error reduction is limited.
@@
See Appendix 4.2 for the proof. From this proposition, we obtain the intuition that the quantization error $\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}$ is bounded by the magnitude of the residual $\left\lVert{\bm{R}}\right\rVert_{F}$. Thus, our goal is to find the optimal ${\bm{L}}_{1}{\bm{L}}_{2}$ that minimizes $\left\lVert{\bm{R}}\right\rVert_{F}=\left\lVert\hat{{\bm{W}}}-{\bm{L}}_{1}{\bm{L}}_{2}\right\rVert_{F}$, which can be solved by Singular Value Decomposition (SVD) (eckart1936approximation; mirsky1960symmetric). Given the SVD of $\hat{{\bm{W}}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}$, the optimal solution is ${\bm{L}}_{1}={\bm{U}}{\bm{\Sigma}}_{:,:r}$ and ${\bm{L}}_{2}={\bm{V}}_{:r,:}$. Figure 5 illustrates the singular value distribution of the original weight $\bm{W}$}, transformed weight $\hat{{\bm{W}}}$ and residual $\bm{R}$}. The singular values of the original weight $\bm{W}$} are highly imbalanced. After smoothing, the singular value distribution of $\hat{{\bm{W}}}$ becomes even sharper, with only the first several values being significantly larger. By removing these dominant values, the magnitude of the residual $\bm{R}$} is dramatically reduced, as $\left\lVert{\bm{R}}\right\rVert_{F}=\sqrt{\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}}$, compared to the original magnitude $\left\lVert\hat{{\bm{W}}}\right\rVert_{F}=\sqrt{\sum_{i=1}^{\min(m,n)}\sigma_{i}^{2}}$, where $\sigma_{i}$ is the $i$-th singular value of $\hat{{\bm{W}}}$. Furthermore, Figure 4(d)(e) show that $\bm{R}$} exhibits fewer outliers with a substantially compressed value range compared to $\hat{{\bm{W}}}$. In practice, we further reduce quantization errors by iteratively updating the low-rank branch through decomposing ${\bm{W}}-Q({\bm{R}})$ and adjusting $\bm{R}$} accordingly for several iterations, and then picking the result with the smallest error.